OpenStudy (anonymous):

A ball is thrown vertically downwards with initial velocity 5 m.s from 30 m above the ground. how long does it take for the ball to hit the ground? g=10 m/s

OpenStudy (abhisar):

Here, initial velocity = 5 m/s Distance = 30 m Acceleration = 10 m/s^2 We need to find time taken, we can use the equation \(\sf s=ut+0.5at^2\) \(\Rightarrow \sf 30=5\times t + 0.5\times 10\times t^2\\ \Rightarrow 5t^2+5t-30=0\)

OpenStudy (abhisar):

ok ?

OpenStudy (abhisar):

Now solving the quadratic equation, we get values of t as -3 and 2

OpenStudy (anonymous):

Okay got it, what formula is this? I'm not quite sure with this formula.

OpenStudy (abhisar):

Since time can not be negative, we take t as 2

OpenStudy (abhisar):

\(\boxed{\sf \huge s=ut+\frac{1}{2}at^2}\)

OpenStudy (anonymous):

I see thanks!

OpenStudy (abhisar):

My pleasure :)